In a circuit , R_1=2.0 Ω,〖 R〗_2=4.0 Ω,R_3=3.0 Ω,〖 R〗_4=6.0 Ω, R_5=12.0 Ω and R_6=2.0 Ω. Calculate the total resistance in the circuit.
1
Expert's answer
2019-07-22T11:29:54-0400

In case of parallel resistors:



"R=\frac{1}{\frac{1}{R_1}+\frac{1}{R_2}+\frac{1}{R_3}+\frac{1}{R_4}+\frac{1}{R_5}+\frac{1}{R_6}}=0.545\space\Omega."

In case of series connection of resistors the total resistance is simply a sum of all resistances, or 29 ohm. It means that series connection give higher resistance and parallel connection provides lower resistance.